Grade 12Analytical Geometry

the equation of directrix and latusectum of a parabola are 3x-4y+27=0 and 3x-4y+2=0 respectively .then the length of latusrectum is???
please give solution...............

Mrunal Sonawane
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear student
 
distance btween equation of directrix and equation of latus rectum = 2a
 
hence 2a = 27 – 2 /(5) = 25/5 = 5
 
since 2a = 5
 
hence 4a = 10
 
means length of latus rectum = 10